Calcium phosphate reacts with sand and carbon to form phosphorus and carbon monoxide and if the reaction is 78.50 %efficient, calculate the mass of phosphorus that maybe be produced from 2.35 kg calcium phosphate, 1.00 kg SiO2 and an excess C. Apart from carbon, what mass of the other excess reactant remained at the end of the reaction?

1
Expert's answer
2021-11-01T06:33:01-0400

Ca3(PO4)2 + 5 C + 3 SiO2 → 2 P + 3 CaSiO3 + 5 CO

n = m / M

M (Ca3(PO4)2) = 310.18 g/mol

M (SiO2) = 60.08 g/mol

M (P) = 30.97 g/mol

M (CaSiO3) = 172.24 g/mol


Available amount of key reactants is:

n (Ca3(PO4)2) = 2 350 / 310.18 = 7.58 mol

n (SiO2) = 1 000 / 60.08 = 16.64 mol


According to the equation, n (P) = 2 x n (Ca3(PO4)2) = 2/3 x n (SiO2). The amount of Phosphorus to be formed from the given amount of key reactants:

n (P) = 2 x n (Ca3(PO4)2) = 2 x 7.58 = 15.16 mol

n (P) = 2/3 x n (SiO2) = 2/3 x 16.64 = 11.1 mol

Therefore, SiO2 is the limiting reactant.


The amount of Phosphorus being formed is: n (P) = 11.1 mol.

m (P) = n x M = 11.1 x 30.97 = 343.77 g

m (P)78.50% = 343.77 x 0.7850 = 269.86 g

n (P)78.50% = 269.86 / 30.97 = 8.71 mol


The remaining amount of Ca3(PO4)2 is:

n (Ca3(PO4)2)excess = n (Ca3(PO4)2) - (n (P) / 2) = 7.58 - (8.71 / 2) = 3.22 mol

m (Ca3(PO4)2)excess = n x M = 3.22 x 310.18 = 998.78 g